The firm is an all-equity firm with assets worth $512,000 and 64,000 shares outstanding. It plans to borrow $120,000 and use these funds to repurchase shares. The firm’s marginal corporate tax is 21%, and it plans to keep its outstanding debt equal to $120,000 permanently. If the firm plan to repurchase shares at $9 per share, what is the expected per share value of equity for the leveraged firm? Please show your work.

A) $8 per share B) $10.45 per share C) $8.23 per share D) $8.59 per share E) Repurchase will not be successful because the firm’s offer is too low.
Answer #1

Current share price = asset worth/ number of share

= 512000/64000

= $8

Price before share repurchase = (512000 + 21%of(120000))/64000

= $537200/64000

= $8.394

With $9 as per share price $120000 will give 13,333.33 shares

Thus remaining share will be equal to 64000 - 13333 = 50667

Share price of levered firm = ((512000 + 21%of(120000)) - 120000)/ 50667

= $8.23 per share

Correct answer is option C.